40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed Aug 4 20:28:03 2010, Alastair, Laser, GYRO, Locking.... 
    Reply  Wed Aug 4 23:19:19 2010, Alastair, Laser, GYRO, Locking.... 
       Reply  Mon Aug 9 18:36:14 2010, rana, Laser, GYRO, Zach is feeeling excluded from this lack of locking IMG_1671.JPG
          Reply  Mon Aug 9 20:48:38 2010, Dmass, Laser, GYRO, Zach is feeeling excluded from this lack of locking 
             Reply  Tue Aug 10 13:44:49 2010, Zach, Laser, GYRO, Zach is feeeling excluded from this lack of locking 
Message ID: 907     Entry time: Wed Aug 4 23:19:19 2010     In reply to: 905     Reply to this: 915
Author: Alastair 
Type: Laser 
Category: GYRO 
Subject: Locking.... 

Quote:

After we took that noise spectrum we started putting in the PD for the transmission readout.  It can only take 55uW according to the spec sheet so we started by turning the laser power down a bit and then we were putting in a BS to reduce the power even more.  We were just getting this set up when the cavity stopped locking properly.  I was working on the theory that it was the lower laser power was causing issues, so I turned it back up - still won't lock.  I checked the beam is on the reflection PDs (it is) and that they weren't saturating (they're not).

We're getting big dips in power on the reflection PD when I scan the cavity through resonance.  We also get a big error signal out of the mixer when I scan the cavity, but it looks kind of one sided (there is a massive positive side and hardly any negative side).  I can also see the error signals from the sidebands resonating, and they look more normal.

What we're seeing when we try to lock is that it only stays locked for a couple of seconds, and we get big fluctuations in the control signal to the piezo, before it falls out of lock.

More investigation will ensue.

PS.  The mojo beads were on the laser - Jenna has removed them in case they were giving it bad vibes.

 Still not got the cavity back to locking properly yet.  It'll have to wait till tomorrow.  I don't understand why the PDH signal has a really large positive side and small negative side when you scan the cavity - it needs more investigating.

I popped out to the 40m around 9pm to check the Rb clocks.  They had come out of lock earlier (possibly due to being too hot inside the igloo).  Rana has put a post up here about them http://nodus.ligo.caltech.edu:8080/40m/3361

They take about 12hrs to lock so the estimate of tomorrow afternoon seems reasonable.

ELOG V3.1.3-